Circular No. 8629

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005ir–2005kc Three apparent supernovae brighter than mag 20 have recently been

reported. SN 2005ir was discovered independently by the Sloan Digital Sky Survey (SDSS) II collaboration (via multiple g, r, and i images taken with the SDSS 2.5-m telescope; communicated by J. Frieman, Fermilab and University of Chicago) and by R. Quimby, M. Sellers, and F. Castro (University of Texas; via unfiltered CCD images with the 0.45-m ROTSE-IIIb telescope at the McDonald Observatory); the tabulated data below are from Quimby et al. (position uncertainty ± 0′′.6). SN 2005kb was also found by the SDSS II group. SN 2005kc was reported by T. Puckett and G. Sostero (cf. IAUC 8615) via unfiltered CCD images with the 0.35-m automated supernova patrol telescope.

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005ir Nov. 3.11 1^{h}16^{m}43.^{s}76 + 0^{o}47′40′′.4 18.5 1′′.7 W, 3′′.4 N
2005kb Nov. 5  0 50 50.68 + 0 51 13.0 18.1 13′′.7 W, 8′′.4 N
2005kc Nov. 9.04 22 34 07.34 + 5 34 06.3 18.2 7′′.6 E, 7′′.4 S

Frieman et al. provide position end figures 43^{s}.80, 40′′.6 for 2005ir; additional approximate magnitudes: Aug. 22 and 28, [18.9 (Quimby et al.); Oct. 28 UT, g = 21.1 (SDSS II); Nov. 6.14, 17.8 (Quimby et al.). Spectroscopy by both groups indicate that 2005ir is a type-Ia supernova with redshift z = 0.08, with peak brightness around now. Additional information on SN 2005ir can be found on CBET 277. SDSS magnitudes of 2005kb: Nov. 5, g = 18.1, r = 18.3, i = 18.5; Nov. 7, 18.0, 18.0, 18.3. A spectrum taken on Nov. 8 with the ARC 3.5-m telescope (+ DIS) shows SN 2005kb to be of type II; an earlier SDSS spectrum shows the host galaxy to be at redshift z = 0.0153 (cf. CBET 281). Additional approximate magnitudes for SN 2005kc in NGC 7311: Sept. 6 and 9, [20.0; Nov. 11.02, 18.0 (0.60-m reflector).

Frieman also reports the discovery of 35 additional supernovae found by the SDSS II collaboration and designated 2005is–2005ka on CBET 280. All were fainter than magnitude g = 20.5 upon discovery in October, and all are type-Ia or probable type-Ia supernovae except for 2005jq and 2005jr (which are of type IIn).

COMET 173P/MUELLER Comet P/1993 W1 = 2005 T1 (cf. IAUC 8613) has been given the

permanent number 173P (cf. MPC 54939).

2005 November 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green